### v1

@@ 3.1 Uniform random field model
$0<a_{\min}:=\|\psi_{0}\|_{\infty}-\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}$ $<a_{\max}:=\|\psi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}<\infty,$
$0<\kappa_{\min}:=\|\xi_{0}\|_{\infty}-\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}$ $<\kappa_{\max}:=\|\xi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}<\infty,$
@@
Recall it is necessary that $\lambda>\kappa_{\max}\geq 0$, which in this case is independent of $\boldsymbol{y}$}. Then, the coefficients in PDE II are
$b^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $:=\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\,e^{\lambda\,t}=\Big(\xi_{0}({\boldsymbol{x}},t)+\sum_{j\geq 1}y_{j}\,\xi_{j}({\boldsymbol{x}},t)\Big)\,e^{\lambda\,t},$ (3.3)
$c^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $:=\lambda+f({\boldsymbol{x}},t)-\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)=\lambda+f({\boldsymbol{x}},t)-\Big(\xi_{0}({\boldsymbol{x}},t)+\sum_{j\geq 1}y_{j}\,\xi_{j}({\boldsymbol{x}},t)\Big).$ (3.4)
@@ 3.2 Parametric regularity for uniform random fields
We will write derivatives with respect to the parameters $\boldsymbol{y}$} using multiindex notation, where for a multiindex $\boldsymbol{\nu}$$\in{\mathbb{N}}^{\mathbb{N}}$ with $|{\boldsymbol{\nu}}|:=\sum_{j\geq 1}\nu_{j}<\infty$, we define the derivative of order $\boldsymbol{\nu}$} by
@@
$\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}\leq 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|},$ (3.5)
where ${\boldsymbol{m}}<{\boldsymbol{\nu}}$ means that $m_{j}\leq\nu_{j}$ for all $j\geq 1$, and $\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}=\prod_{j\geq 1}\binom{\nu_{j}}{m_{j}}$ is a product of binomial coefficients.
Theorem 3.1 (parametric regularity)
For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5). For $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, let $w^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the corresponding solution to PDE II (2.8)–(2.10). Then for every multiindex $\boldsymbol{\nu}$}, the $\boldsymbol{\nu}$}-th derivative with respect to $\boldsymbol{y}$} of the solutions satisfy the regularity bounds
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}w^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\max(C,\tfrac{E}{\rho})\,(\rho\,{\boldsymbol{\beta}})^{\boldsymbol{\nu}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)},$
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}u^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\widetilde{C}\,(\rho\,{\boldsymbol{\beta}})^{\boldsymbol{\nu}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)},\quad\widetilde{C}:=\sqrt{\max(1+2\lambda^{2},2)}\,e^{\lambda\,T}\,\max(C,\tfrac{E}{\rho}),$
@@
where $C$ is defined in Theorem 2.2, while $E$ and $\rho$ are defined in (3.20) and (3.21) below, respectively, both depending on $\Phi$ defined in (3.18) below.
For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$, we have the mixed derivative $\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}$ of (3.1)–(3.4)
$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}a^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}\psi_{j}({\boldsymbol{x}},t)\hskip 14.22636pt&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise},\end{cases}$ (3.6)
$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}b^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}\xi_{j}({\boldsymbol{x}},t)\,e^{\lambda\,t}&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise},\end{cases}$ (3.7)
$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}c^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}-\xi_{j}({\boldsymbol{x}},t)\hskip 8.5359pt&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise}.\end{cases}$ (3.8)
Recall that the weak solution $w\in{\mathcal{X}}$ satisfies (2.12). For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ we take the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the first condition in (2.12) and use the Leibniz product rule to obtain for all $v\in{\mathcal{X}}$,
$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $=-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{{\boldsymbol{k}}}a)\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}b)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w^{2})\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}c)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t,$ (3.9)
where the mixed derivative of $w^{2}$ can be expanded further using the Leibniz product rule as
$(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w^{2})$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{k}}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{k}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}-{\boldsymbol{m}}}w),$
@@
$(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}}w^{2})$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)=2\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w+\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w).$
Separating out the ${\boldsymbol{k}}={\boldsymbol{0}}$ terms in (3.2), applying (3.6)–(3.8), and rearranging to have the highest-order mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w$ on the left-hand side, we obtain
$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad+\int_{I}\int_{\Omega}a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
@@
$=-\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.10)
Gradient. Taking the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the second condition in (2.12) and then taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,0)$, we conclude that $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$. This yields
$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}.$
Taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w$ in (3.2) then gives
$\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}+\int_{I}\int_{\Omega}a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
@@
$=-\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$
Until this point we still have equality in (3.2). For the special cases $|{\boldsymbol{\nu}}|=1$, the sum over ${\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}$ is empty and takes the value 0, with no meaning attached to the summand.
We obtain a lower bound on the left-hand side of (3.2) by dropping some nonnegative terms, including the term with integrand $2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,w\geq 0$ since $0\leq w\leq e^{-\lambda\,t}$, so

### v2

@@ 3.1 Uniform random field model
$0<a_{\min}:=\min_{t\in I}\inf_{{\boldsymbol{x}}\in\Omega}|\psi_{0}({\boldsymbol{x}},t)|-\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}<a_{\max}:=\|\psi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}<\infty,$
$0\leq\kappa_{\min}:=\min_{t\in I}\inf_{{\boldsymbol{x}}\in\Omega}|\xi_{0}({\boldsymbol{x}},t)|-\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}<\kappa_{\max}:=\|\xi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}<\infty,$
@@ 3.2 Parametric regularity for uniform random fields
We will write derivatives with respect to the parameters $\boldsymbol{y}$} using multiindex notation, where for a multiindex ${\boldsymbol{\nu}}\in{\mathbb{N}}^{\mathbb{N}}_{0}$ with $|{\boldsymbol{\nu}}|:=\sum_{j\geq 1}\nu_{j}<\infty$, we define the derivative of order $\boldsymbol{\nu}$} by
@@
$\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}\leq 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|},$ (3.3)
where ${\boldsymbol{m}}<{\boldsymbol{\nu}}$ means that ${\boldsymbol{m}}\neq{\boldsymbol{\nu}}$ and $m_{j}\leq\nu_{j}$ for all $j\geq 1$, and $\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}=\prod_{j\geq 1}\binom{\nu_{j}}{m_{j}}$ is a product of binomial coefficients.
Theorem 3.1 (parametric regularity for uniform model)
For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5) defined by affine expansions (3.1) and (3.2). Then for every multiindex $\boldsymbol{\nu}$}, the $\boldsymbol{\nu}$}-th derivative with respect to $\boldsymbol{y}$} of the solution satisfies $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u^{\boldsymbol{y}}\in{\mathcal{X}}$ with the regularity bound
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}u^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq\widetilde{C}\,(\rho\,{\boldsymbol{\beta}})^{\boldsymbol{\nu}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)},\quad\widetilde{C}:=\max(C,\tfrac{E}{\rho}),$
@@
where $C$ is defined in Theorem 2.2 (in the uniform model, $a_{\max}$, $a_{\min}$, $\kappa_{\max}$, $f_{\max}$ are independent of $\boldsymbol{y}$}), while $E$ and $\rho$ are defined in (3.22) and (3.23) below, respectively, both depending on $\Phi$ defined in (3.20) below.
For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$, we have the mixed derivative $\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}$ of (3.1) and (3.2)
$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}a^{\boldsymbol{y}}({\boldsymbol{x}},t)=\begin{cases}\psi_{j}({\boldsymbol{x}},t)&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$},\\ 0&\mbox{otherwise},\end{cases}\quad\mbox{and}\quad\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)=\begin{cases}\xi_{j}({\boldsymbol{x}},t)&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$},\\ 0&\mbox{otherwise}.\end{cases}$ (3.4)
Recall that the weak solution $u\in{\mathcal{X}}$ satisfies (2.12). For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ we take the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the first condition in (2.12), but we integrate over $t\in[0,\tau]$ for some $\tau\leq T$ instead of the full interval $I$. Using the Leibniz product rule, the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ for the space-time integrand in (2.12) can be written as
$\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}\big[a\,\nabla u\cdot\nabla v+(\kappa\,u^{2}+(f-\kappa)\,u)\,v\big]$
$=\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}\Big[(\partial_{\boldsymbol{y}}^{{\boldsymbol{k}}}a)\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}u)\cdot\nabla v+(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}\kappa)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}u^{2})\,v+(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}(f-\kappa))\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}u)\,v\Big],$ (3.5)
where the mixed derivative of $u^{2}$ can be expanded further using the Leibniz product rule as
$\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}u^{2}$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{k}}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{k}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}-{\boldsymbol{m}}}u),$
@@
$\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}}u^{2}$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)=2\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,u+\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u).$
Separating out the ${\boldsymbol{k}}={\boldsymbol{0}}$ terms in (3.2) and applying (3.4), we obtain for all $v\in{\mathcal{X}}$,
@@
$\int_{0}^{\tau}\!\!\big\langle\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u),v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\bigg[a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\cdot\nabla v+2\,\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,u\,v+(f-\kappa)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,v$ (3.6)
$\qquad+\sum_{j\geq 1}\nu_{j}\,\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\cdot\nabla v-\sum_{j\geq 1}\nu_{j}\,\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\,v+\kappa\,\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)\,v$
$\qquad+\sum_{j\geq 1}\nu_{j}\,\xi_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u)\,v\bigg]\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0.$
Membership in $\mathcal{X}$}. Taking $\tau=T$ in (3.6) and moving all terms involving derivatives of order strictly less than $\boldsymbol{\nu}$} to the right-hand side, we can see that $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u$ satisfies (the weak form of) a linear parabolic PDE
@@
$\int_{I}\big\langle\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u),v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}\big[a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\cdot\nabla v+\widetilde{b}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,v\big]\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ (3.7)
$=\int_{I}\langle F_{\boldsymbol{\nu}},v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t,\qquad\forall v\in{\mathcal{X}},$
where $\widetilde{b}=\widetilde{b}^{\boldsymbol{y}}({\boldsymbol{x}},t)=2\,\kappa\,u+f-\kappa\in L^{2}(I;L^{\infty}(\Omega))$ and $F_{\boldsymbol{\nu}}=F_{\boldsymbol{\nu}}(t)$ is the (time-dependent) functional defined by